OpenStudy (bloomlocke367):

I need help determining whether a function is even or odd. I know HOW to...for the most part, I just need a little bit of clarification.

OpenStudy (bloomlocke367):

For \(g(x)=x^3\) you have to plug in -x and see if the function is the same or completely the opposite. I know if it's the same that the function is even, and if it's completely the opposite it's odd. BUT my teacher did a few examples on the board where she took out -1, so the function would be even when you'd think it would be odd. When would you take out a -1?

OpenStudy (bloomlocke367):

Sorry for making it so wordy XD

Nnesha (nnesha):

you would multiply by -1 to check if it's odd or neither

OpenStudy (danjs):

Even Function is symmetric about Y axis F(x) = F(-x)

OpenStudy (bloomlocke367):

yes @DanJS. That's what I learned.

OpenStudy (danjs):

the odd is reflected over the origin, like the line y=x , or y = x^3

OpenStudy (bloomlocke367):

and for odd you have f(-x)=-f(x)

OpenStudy (bloomlocke367):

ohh okay. I'm still confused on what she did in class today XD

OpenStudy (danjs):

yep, that makes sense

Nnesha (nnesha):

for exxample if f(X)=x^3-x i would substitute x for -x (-x)^2 -(-x) -x^3 +x <-- not the original equation so it's not even now i would multiply by -1 (-x^3 -x) times -1 = x^3-x

OpenStudy (bloomlocke367):

Now you're confusing me :O

OpenStudy (danjs):

Even functions are symmetric about the y-axis, Odd functions are symmetric about both x and y-axis, that's it

OpenStudy (bloomlocke367):

odd are symmetric about both? so about the origin...or...?

OpenStudy (danjs):

right, about the origin, same as reflected over both axis

OpenStudy (bloomlocke367):

okay thanks XD so \(f(x)=\sqrt{x^2+2}\) is even, right? because when you plug in -x the equation stays the same and when you graph it it's a porabola.

OpenStudy (danjs):

Just switch x for -x and see if you get either the original f(x) - or the negative - f(x) or neither

OpenStudy (danjs):

that is basically all you need to do

OpenStudy (bloomlocke367):

yes, I know. you get \(f(-x)=\sqrt{(-x)^2+2}\) which simplifies to \(f(-x)=\sqrt{x^2+2}\) which is the same as the original, right?

OpenStudy (danjs):

yea you got it...but it is not parabola, y^2 - x^2 = k hyperbola, upper half +root

OpenStudy (bloomlocke367):

oops XD I only saw the upper half. My bad. Thanks.

OpenStudy (danjs):

it is only the upper half, not + or - root, parabola is y = x^2 no square on the y.. eh nevermind

OpenStudy (bloomlocke367):

And I also got even for \(\large g(x)=\frac{3}{1+x^2}\) and yes, I should have realized. I did study hyperbolas and conics last year in Alg 2

OpenStudy (danjs):

yeah , if the only x term is squared, it will be even prolly

OpenStudy (danjs):

you get it all good now?

OpenStudy (bloomlocke367):

Yes, I THINK so. I'll tag you if I have any more questions.

OpenStudy (danjs):

just change to -x if you get f(x) back ---even if you get -f(x) back ---odd --------------------------- most functions are neither even or odd

OpenStudy (bloomlocke367):

WAIT. I think this is neither. \(f(x)=-x^2+0.03x+5\)

OpenStudy (bloomlocke367):

andddd he left.

OpenStudy (bloomlocke367):

@DanJS

OpenStudy (bloomlocke367):

@Nnesha do you know that one?

OpenStudy (bloomlocke367):

http://prntscr.com/8bwwtw there's the graph.

OpenStudy (bloomlocke367):

it's symmetrical over the y-axis but when you plug in -x it doesn't work out so I'm confused.

Nnesha (nnesha):

\(\huge\color{green}{\checkmark}\)

Nnesha (nnesha):

well i don't know abt graph but when u replace x with -x you will get -x^2-.03x+5 right ? so it's not even now i would multiply by -1 \[-1(-x^2-0.03x+5)=x^2+0.03-5\] i didn't get the original equation -x^2+0.03x+5 that's how its neither and ~~this is how we di it!

OpenStudy (mathmate):

Some simple rules, Sum of odd functions is odd. Sum of even functions is even. Sum of odd and even is neither!

OpenStudy (bloomlocke367):

What? What do you mean mm?

OpenStudy (mathmate):

and a constant is even.

Nnesha (nnesha):

did*:XD

OpenStudy (mathmate):

f(x)=x^2+x is a sum of even and odd functions, so f(x) is neither odd nor even!

OpenStudy (bloomlocke367):

I really don't understand what you mean mm XD

OpenStudy (mathmate):

If you are given a function which is the sum of many terms,

OpenStudy (mathmate):

Check out each term individually.

OpenStudy (mathmate):

Check if "each" term is even or odd.

OpenStudy (mathmate):

For example, f(x)=x^4+3x^2 is even because x^4 and 3x^2 are both even.

OpenStudy (bloomlocke367):

ohhhh. That's not what I learned but okay. XD I just plug -x where every x is and solve to see if the equation matches the original or not.

OpenStudy (bloomlocke367):

but your method seems easier XD Thank you.

OpenStudy (mathmate):

No, this does not replace what you learned. This makes your life easier when you have a function consisting of multiple terms. I assume you already know that x^2, x^4... are even, and x, x^3.... are odd.

OpenStudy (bloomlocke367):

yes. and I know, I have to prove these problems algebraically and graphically anyways. But it's good to know for the future.

OpenStudy (mathmate):

Good! Are you all clear now, for example, is y=x+3 even or odd?

OpenStudy (bloomlocke367):

so \(g(x)=2x^3-3x\) is odd, right? because after you plug in -x you get \(g(-x)=-2x^3+3x\)

OpenStudy (bloomlocke367):

and that's neither... right? The one you showed?

OpenStudy (mathmate):

exactly!

OpenStudy (mathmate):

Exactly again!

OpenStudy (bloomlocke367):

Awesome, thank you!

OpenStudy (mathmate):

Great! You're welcome! :)

OpenStudy (bloomlocke367):

woe woe woe. I have one more. This one is like the example she gave us.

OpenStudy (bloomlocke367):

\(\large h(x)=\frac{1}{x}\)

OpenStudy (mathmate):

Apply the rules and see what you get! :)

OpenStudy (bloomlocke367):

you get \(h(-x)=\large\frac{1}{-x}\) but you can take \(h(-x)=\large\frac{1}{-1(x)}\) which would make it even.

OpenStudy (mathmate):

given h(x)=1/x h(-x)=1/(-x)=-1/x=-h(x) so...

OpenStudy (bloomlocke367):

what? I'm so confused

OpenStudy (bloomlocke367):

so it's odd?

OpenStudy (bloomlocke367):

because -h(x)

OpenStudy (mathmate):

After you have put in the -x, you want to rearrange the expression in terms of h(x) so you can decide if h(x) is even or odd. You have these rules: if h(-x)=-h(x), then h(x) is odd if h(-x)=h(x), then h(x) is even.

OpenStudy (mathmate):

yes, it is odd. It's easier to visualize it graphically.

OpenStudy (bloomlocke367):

so it's odd then. because it would equal \(-\frac{1}{x}\)?

OpenStudy (bloomlocke367):

and this is the graph. http://prntscr.com/8bx7i1

OpenStudy (mathmate):